# A new infinitesimal form of the Prékopa–Leindler inequality with multiplicative structure and applications

### v1

@@ 5 Symmetric case and Hessian pinching
In this final section, we shall prove Theorem 1.4. To do so, we specialize to the case where the convex body $K$ is origin-symmetric and the measure $\mu$ is even. We begin with the following lemma.
Lemma 5.1.
$k_{1}\,\mathrm{Id}\leq\nabla^{2}u.$
Then, if $\mu$ is such that $\mathrm{d}\mu(x)=e^{-u(x)}\,\mathrm{d}x$,
$\frac{1}{\mu(K)}\int_{K}\|x\|^{2}\,\,\mathrm{d}\mu\leq\frac{n}{k_{1}}.$ (69)
Proof.
Since $K$ is origin-symmetric and $\mu$ is even, for each $i\in\{1,\dots,n\}$ we have $\int_{K}x_{i}\,\,\mathrm{d}\mu=0$. By the Brascamp–Lieb inequality,
$\int_{K}x_{i}^{2}\,\,\mathrm{d}\mu\leq\int_{K}\langle(\nabla^{2}u)^{-1}e_{i},e_{i}\rangle\,\,\mathrm{d}\mu,$
where $e_{i}$ denotes the $i$-th coordinate vector. Summing over $i$ yields
$\int_{K}\|x\|^{2}\,\,\mathrm{d}\mu\leq\int_{K}\mathrm{tr}\big((\nabla^{2}u)^{-1}\big)\,\,\mathrm{d}\mu\leq\frac{n}{k_{1}}\,\mu(K),$
proving the claim. ∎
We shall also use the following lemma, borrowed from [23]; a proof can be given along the same lines as Lemma 5.1.
@@
$\frac{1}{p(\mu,K)}\leq 2n+\frac{1}{\mu(K)}\int_{K}\Big(\langle(\nabla^{2}u)x,x\rangle+\langle(\nabla^{2}u)^{-1}\nabla u,\nabla u\rangle\Big)\,\,\mathrm{d}\mu.$
@@
Let $g(x)=\langle\nabla u(x),x\rangle$. By (68) and choosing $\rho=-\overline{\rho}$ and $\varphi=g$ in Theorem 2.6
$\int_{\partial K}h_{K}(\nu_{K})\,\,\mathrm{d}\mu-\int_{\partial K}\overline{\rho}\,\,\mathrm{d}\mu=\langle\overline{\rho},g\rangle_{\rm I}\geq-\frac{\langle g,g\rangle_{\rm BL}+\langle\overline{\rho},\overline{\rho}\rangle_{\rm P}}{2}.$
Using $\langle\overline{\rho},\overline{\rho}\rangle_{\rm P}=\int_{\partial K}\overline{\rho}\,\,\mathrm{d}\mu$, this implies
$\frac{1}{\mu(K)}\int_{\partial K}\overline{\rho}\,\mathrm{d}\mu\leq\frac{\langle g,g\rangle_{\rm BL}}{\mu(K)}+\frac{2}{\mu(K)}\int_{\partial K}h_{K}(\nu_{K})\,\mathrm{d}\mu.$ (71)
Direct computations dropping the non-positive variance term show that
$\langle g,g\rangle_{\rm BL}\leq\int_{K}2\langle\nabla u,x\rangle+\langle(\nabla^{2}u)x,x\rangle+\langle(\nabla^{2}u)^{-1}\nabla u,\nabla u\rangle\,\mathrm{d}\mu,$
and the proof is concluded with an application of (65) upon noting that $\frac{1}{p(\mu,K)}=\frac{1}{\mu(K)}\int_{\partial K}\overline{\rho}$. ∎
The above estimate cannot lead to a resolution of the dimensional Brunn–Minkowski conjecture, since even applying the main inequality from Theorem 1.1 already yields a bound strictly larger than $n$. Nevertheless, the estimate holds for an arbitrary convex body $K$, which need not contain the origin, and for an arbitrary smooth log-concave measure $\mu=e^{-u}$. In this general setting, we obtain control of the concavity power in terms of the quantity
$\langle(\nabla^{2}u)x,x\rangle+\langle(\nabla^{2}u)^{-1}\nabla u,\nabla u\rangle.$
Assuming in addition that $K$ is symmetric and that the Hessian of $u$ is pinched, one can still deduce bounds linear in the dimension, in the spirit of [23].
@@
$k_{1}\,\mathrm{Id}\leq\nabla^{2}u\,\,\text{and}\,\,\,\Delta u\leq k_{2}n.$
If $\mu$ is such that $\mathrm{d}\mu(x)=e^{-u(x)}\,\mathrm{d}x$, setting $r=\frac{k_{2}}{k_{1}}$ we have
$p(\mu,K)\geq\frac{c}{n},$ (72)
with $c=\frac{1}{2(r+1)}$. In particular,
@@
By Proposition 5.3 together with Lemma 5.2, we obtain
$\frac{1}{p(\mu,K)}\leq 2n+rn+\frac{1}{\mu(K)}\int_{K}\langle(\nabla^{2}u)^{-1}x,x\rangle\,\,\mathrm{d}\mu.$
Since $(\nabla^{2}u)^{-1}\leq k_{2}\,\mathrm{Id}$ and by (69),
$\frac{1}{\mu(K)}\int_{K}\langle(\nabla^{2}u)^{-1}x,x\rangle\,\,\mathrm{d}\mu\leq k_{2}\,\frac{n}{k_{1}}=rn.$
Combining the above estimates yields
$\frac{1}{p(\mu,K)}\leq(2r+2)n,$
which is equivalent to (72). ∎

### v2

@@ 5 Symmetric case and Hessian pinching
In this final section, we shall prove Theorem 1.4. To do so, we specialize to the case where the convex body $K$ is origin-symmetric and the measure $\mu$ is even. First we report the following lemma, which is a direct consequence of [23, Lemma 5.1].
@@
$\frac{1}{p(\mu,K)}\leq n+\frac{1}{\mu(K)}\int_{K}\langle(\nabla^{2}u)^{-1}\nabla u,\nabla u\rangle\,\,\mathrm{d}\mu.$
@@
Let $g(x)=\langle\nabla u(x),x\rangle$ and $f(x)=\overline{\rho}(x)-h_{K}(\nu_{K}(x))$. By (69) we infer that
$\mathcal{L}(f)=\mathcal{L}(\overline{\rho}-h_{K}(\nu_{K}))=-\langle\nabla u,x\rangle+\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\mu.$
Multiplying both sides by $f$ and integrating on $\partial K$ with respect to $\mu$, by (64) we deduce $\langle f,f\rangle_{{\rm P}}=-\langle g,f\rangle_{\rm I}$. Now, the choice $\rho=f$ and $\varphi=-g$ in Theorem 2.6 yields
$\langle f,f\rangle_{{\rm P}}=-\langle g,f\rangle_{\rm I}\geq-\frac{\langle f,f\rangle_{{\rm P}}+\langle g,g\rangle_{{\rm BL}}}{2}\implies-\langle f,f\rangle_{{\rm P}}=\langle g,f\rangle_{\rm I}\geq-\langle g,g\rangle_{{\rm BL}}.$ (71)
Furthermore, notice that through an integration by parts and using (66)
$\langle f,h_{K}(\nu_{K})\rangle_{{\rm I}}$ $=\int_{\partial K}\langle\nabla u,x\rangle h_{K}(\nu_{K})\,\mathrm{d}\mu-\frac{1}{\mu(K)}\int_{\partial K}h_{K}(\nu_{K})\,\mathrm{d}\mu\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\mu$
$=n\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\mu-\int_{K}\langle\nabla u,x\rangle^{2}\,\mathrm{d}\mu+\int_{K}\langle(\nabla^{2}u)x,x\rangle\,\mathrm{d}\mu+\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\mu$
$-n\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\mu+\frac{1}{\mu(K)}\left(\int_{K}\langle\nabla u,x\rangle\right)^{2}\,\mathrm{d}\mu$
$=\int_{K}\langle(\nabla^{2}u)x,x\rangle\,\mathrm{d}\mu+\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\mu-\int_{K}\langle\nabla u,x\rangle^{2}\,\mathrm{d}\mu+\frac{1}{\mu(K)}\left(\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\mu\right)^{2}.$ (72)
Combining (71) and(5) with (69), we infer
$n\mu(K)-\int_{\partial K}\overline{\rho}\,\mathrm{d}\mu$ $=\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\mu+\langle g,f\rangle_{{\rm I}}+\langle g,h_{K}\rangle_{{\rm I}}\geq\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\mu-\langle g,g\rangle_{{\rm BL}}+\langle g,h_{K}\rangle_{{\rm I}}$
$=\int_{K}\langle(\nabla^{2}u)x,x\rangle\,\mathrm{d}\mu+2\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\mu-\int_{K}\langle\nabla u,x\rangle^{2}\,\mathrm{d}\mu+\frac{1}{\mu(K)}\left(\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\mu\right)^{2}$
$-\int_{K}\langle(\nabla^{2}u)x,x\rangle\,\mathrm{d}\mu-2\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\mu+\int_{K}\langle\nabla u,x\rangle^{2}\,\mathrm{d}\mu-\frac{1}{\mu(K)}\left(\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\mu\right)^{2}$
$-\int_{K}\langle(\nabla^{2}u)^{-1}\nabla u,\nabla u\rangle\,\mathrm{d}\mu.$
That is,
$\int_{\partial K}\overline{\rho}\,\mathrm{d}\mu\leq n\mu(K)+\int_{K}\langle(\nabla^{2}u)^{-1}\nabla u,\nabla u\rangle\,\mathrm{d}\mu,$
and the proof is concluded by recalling that $\frac{1}{p(\mu,K)}=\frac{1}{\mu(K)}\int_{\partial K}\overline{\rho}\,\mathrm{d}\mu$. ∎
We remark that the above estimate cannot lead to a resolution of the dimensional Brunn–Minkowski conjecture, since even applying the main inequality from Theorem 1.1 already yields a bound strictly larger than $n$. Nevertheless, the estimate holds for an arbitrary convex body $K$, which need not contain the origin, and for an arbitrary smooth log-concave measure $\mu=e^{-u}$. In this general setting, we obtain control of the concavity power in terms of the quantity
$\frac{1}{\mu(K)}\int_{K}(\nabla^{2}u)^{-1}\nabla u,\nabla u\rangle\,\mathrm{d}\mu.$
Assuming, in addition, that $K$ is symmetric and that the Hessian of $u$ is pinched, we can finally prove Theorem 1.4. We repeat the statement for the convenience of the reader.
@@
Consider an origin-symmetric smooth convex body $K\subset\mathbb{R}^{n}$ and an even strictly convex function $u\in C^{2}(\mathbb{R}^{n})$ such that
$k_{1}\,\mathrm{Id}\leq\nabla^{2}u\,\,\text{ and }\,\,\Delta u\leq k_{2}\,n$
for some $0<k_{1}\leq k_{2}$. If $\mu$ is such that $\mathrm{d}\mu(x)=e^{-u(x)}\,\mathrm{d}x$, setting $r=\frac{k_{2}}{k_{1}}$ and $c=\frac{1}{r+1}$ we have
$p(\mu,K)\geq\frac{c}{n}.$ (73)
In particular,
@@
The proof is a straightforward application of Proposition 5.2 together with Lemma 5.1. ∎